Tom and Jerry's has 2.0 million shares of common stock outstanding, 2.0 million shares of preferred stock outstanding, and 10.00 thousand bonds. If the common shares are selling for $13.00 per share, the preferred share are selling for $10.00 per share, and the bonds are selling for 100.00 percent of par, what would be the weight used for equity in the computation of Tom and Jerry's WACC?
Answer #1

Total value of:

Common stock=(2*13)=$26 million

Preferred stock=(2*10)=$20 million

Bonds=10,000*(1000*100%)=$10 million

Total value=(26+20+10)=$56 million

Hence weight for equity =(26/56)

=46.43%(Approx).